Calcuate the amount of change that passes a cross section each second and the current density of a current of 5.0 a exists in a wire of 0.50 cm diometers
Answer #1

Sol::

Given::

Current(I)= 5A
Diameter(D)= 0.5 cm
Radius(r)= 0.25 cm
r = 0.0025 m

A)

The amount of charge that passes a cross section in a second is current

So

I= Q/t

Where t= 1s

So

Q = I*t
= 5*1
= 5 C

B)

Current density will be
J = I/A
= I/(πr²)
= 5/(π*0.0025²)
= 254695.956 A/m²
≈ 2.54*10^5 A/m²

Hope this answer helps you
Hope will get the VOTE
Thank you